# Defend Data Poisoning Attacks on Voice Authentication

### v1

@@ 7 Conclusion
In this paper, we investigate a targeted data poisoning attack that allows the attacker to impersonate a legitimate user via voice authentication. We propose a novel CNN-based discriminator called Guardian to help distinguish the attacked accounts from normal accounts. We design a series of advanced techniques for the Guardian network to obtain balanced training samples and augmented input feature vectors, which significantly improves the Guardian network’s effectiveness. Our experimental results demonstrate that the Guardian network achieves around 95% detection accuracy while existing defense mechanisms only yield 60% accuracy.

### v2

@@ 7 Conclusion
In this paper, we investigate a targeted data poisoning attack that allows the attacker to impersonate a legitimate user via voice authentication. We propose a novel CNN-based discriminator called Guardian to help distinguish the attacked accounts from normal accounts. We design a series of advanced techniques for the Guardian network to obtain balanced training samples and augmented input feature vectors, which significantly improves the Guardian network’s effectiveness. Our experimental results demonstrate that the Guardian network achieves around 95% detection accuracy while existing defense mechanisms only yield 60% accuracy. In the future, we are interested in exploring new attack scenarios and the effectiveness of our Guardian model against non CNN-based voice authentication models.